### 19 Apr 2026 v1

# Alternating and symmetric separability of free products

Abstract.
Let $F\ast G$ be a free product of a free group $F$ and a LERF group $G$. In this note, we provide sufficient conditions for a subgroup $H$ of $F\ast G$ to be $\mathcal{A}\cup\mathcal{S}$-separable, that is, for any finite set $\{\gamma_{1},\ldots,\gamma_{n}\}\subset(F\ast G)\setminus H$, there is a surjection $f$ from $F\ast G$ to an alternating or symmetric group such that $f(\gamma_{i})\notin f(H)$ for all ii. As a corollary, any finitely generated infinite-index subgroup of a free group is $\mathcal{A}\cup\mathcal{S}$-separable in the free product of the free group and an arbitrary LERF group, generalizing a result of Wilton.
Key words and phrases:
Locally extended residually finite, alternating and symmetric separable, free product, precover, relative Cayley graph
2020 Mathematics Subject Classification:
20F65, 20E26, 57M07.
The authors are partially supported by NSFC (No. 12471066) and the Shaanxi Fundamental Science Research Project for Mathematics and Physics (No. 23JSY027).
1. Introduction
Residual properties have been widely discussed on the topic of infinite discrete groups. Let $\mathcal{C}$ be a class of groups. A group $G$ is said to be residually $\mathcal{C}$ if for every $g\in G\setminus\{e\}$, there exists a homomorphism $f:G\to C$ onto some group $C\in\mathcal{C}$, such that $f(g)$ is nontrivial. The property of residual finiteness admits several generalizations, most notably the concept of LERF: a group $G$ is said to be locally extended residually finite (LERF) if for any finitely generated subgroup $H\leq G$ and any $\gamma\in G\setminus H$, there is a homomorphism $f$ from $G$ onto a finite group such that $f(\gamma)\notin f(H)$.
The LERFness of groups is a property closely related to low dimensional topology and has been widely studied. Large classes of groups are known to be LERF: for example, finitely generated free groups [6], closed orientable hyperbolic surface groups [13], compact Seifert fiber space groups [13], Sol manifold groups [8] and hyperbolic 3-manifold groups [1] [19]. In fact, by [15], a compact (with empty or tori boundary) orientable irreducible 3-manifold $M$ with $\pi_{1}(M)$ LERF if and only if $M$ supports one of Thurston’s eight geometries. Moreover, by [17], limit groups are LERF, which generalizes both the cases of free groups and surface groups. However, nontrivial graph manifold groups are not LERF [11], noncompact arithmetic hyperbolic $m$-manifold groups with $m>3$ and compact arithmetic hyperbolic $m$-manifold groups with $m>4$ (except when defined by octonions with $m=7$) are not LERF [15].
The LERFness admits a broader generalization through the restriction of its finite quotients to a more specific class, for example, the class of alternating groups $\mathcal{A}$ and symmetric groups $\mathcal{S}$.
Definition $1.$
Let $H$ be a subgroup of a finitely generated group $G$ and let $\mathcal{C}$ be a class of groups. We say that $H$ is $\mathcal{C}$-separable in $G$ if for any finite set $\{\gamma_{1},\ldots,\gamma_{n}\}\subset G\setminus H$, there is a surjective homomorphism $f$ from $G$ to a group in $\mathcal{C}$ such that $f(\gamma_{i})\notin f(H)$ for all ii.
Note that $\mathcal{C}$-separability is a property of a certain subgroup, in particular, the trivial subgroup of $G$ is $\mathcal{C}$-separable implies that $G$ is residually $\mathcal{C}$.
Notations. Throughout this paper, let $F_{r}$ be a free group of rank $r>1$ and let $H^{g}=gHg^{-1}$ be the conjugation of a subgroup $H$ by $g$. For a set $V$, let jVj|V| denote the cardinality of $V$.
In [18], Wilton showed that the finitely generated infinite-index subgroups of free groups are $\mathcal{A}$-separable. More precisely,
Theorem 1.2.
[18, Theorem A] Let $H$ be a finitely generated subgroup of infinite index in the free group $F_{r}$ and let $\{\gamma_{1},\ldots,\gamma_{n}\}$ be any finite set of $F\setminus H$. Then there is a surjection $f$ from $F_{r}$ to some finite alternating group $A_{k}$ such that $f(\gamma_{i})\notin f(H)$ for all ii.
Wilton also conjectured that the same result of Theorem 1.2 holds for closed orientable hyperbolic surface groups. In 2021, Buran [2] confirmed Wilton’s conjecture. Moreover, he generalized this result to right-angled Coxeter groups and provided a complete characterization of the $\mathcal{A}$-separability and the $\mathcal{S}$-separability of convex-cocompact subgroups.
We now investigate the residual properties and the $\mathcal{C}$-separability of free products of groups. It is well known that both of the residually finiteness and LERFness are preserved under taking free products, see [12] and [3]. In 1984, Tamburini and Wilson [16] studied the residually alternating property of the free products of residually finite groups.
Theorem 1.3.
[16, Theorem 1] Let $G$ be a free product of at least two nontrivial residually finite groups. Then $G$ is residually alternating unless $G$ is the infinite dihedral group $\mathbb{Z}/2\mathbb{Z}\ast\mathbb{Z}/2\mathbb{Z}$.
Note that LERF groups are residually finite, and the trivial subgroup of $G$ is $\mathcal{A}$-separable implies that $G$ is residually alternating. It is natural to ask the following question.
Question 1.4.
Which kinds of subgroups are $\mathcal{A}$-separable or $\mathcal{S}$-separable in free products of LERF groups?
By combining the methods in [5, 18], we show that there exists a class of $\mathcal{A}\cup\mathcal{S}$-separable subgroups in the free product of a LERF group with a free group $F_{r}$ of rank $r>1$.
Theorem 1.5.
Let $G$ be a LERF group and let $H\leq F_{r}\ast G$ be a finitely generated subgroup. Then $H$ is $\mathcal{A}\cup\mathcal{S}$-separable in $F_{r}\ast G$ if one of the following holds:
  1. (1)
    $H\cap F_{r}^{\gamma}=1$ for every $\gamma\in F_{r}\ast G$, or
  2. (2)
    $H\cap F_{r}^{\gamma}\neq 1$ has infinite index in $F_{r}^{\gamma}$ for some $\gamma\in F_{r}\ast G$.
Since every finitely generated infinite-index subgroup of $F_{r}$ satisfies the hypothesis in Theorem 1.5, the following immediate corollary represents a generalization of Theorem 1.2 regarding the $\mathcal{A}$-separability of free groups.
Corollary 1.6.
Any finitely generated infinite-index subgroup of $F_{r}$ is $\mathcal{A}\cup\mathcal{S}$-separable in the free product $F_{r}\ast G$, where $G$ is a LERF group.
Notice that $\mathcal{A}\cup\mathcal{S}$-separable means the quotient is required to be either an alternating or a symmetric group. Since $G$ is an arbitrary LERF group, no uniform modification method seems to exist that would guarantee the quotient is precisely an alternating group. Moreover, the hypothesis “infinite index" is essential. In fact, the following example shows there exist finite-index subgroups that are not alternating separable in such a free product.
Example 1.7.
[…]
$\phi:F_{2}\ast G\to\mathbb{Z}/2\mathbb{Z}$
given by
$\phi(a)=\phi(b)=1,\quad\phi(c)=0.$
Then $\ker\phi$ is a finitely generated normal subgroup of $F_{2}\ast G$ of index 2. For any surjection $f:F_{2}\ast G\to A_{n}$, $f(\ker\phi)$ is normal in $A_{n}$ with index at most 2. If $n=3$ or $n\geq 5$, then $A_{n}$ is simple and hence $f(\ker\phi)=A_{n}$; if $n=4$, then $f(\ker\phi)$ is again $A_{n}$ because the only proper normal subgroup of $A_{4}$ is of index 3. So $\ker\phi$ is not $\mathcal{A}$-separable.
To prove Theorem 1.5, we will construct a relative Cayley graph of $F_{r}\ast G$. This graph corresponds to the 1-skeleton of a primary sheeted cover of the representation complex associated with $F_{r}\ast G$. The action of $F_{r}\ast G$ on its vertex set will then be used to demonstrate $\mathcal{A}\cup\mathcal{S}$-separability.
This paper is organized as follows. In Section 2, we review some topological tools, such as labeled graphs and precovers, to prepare for proving the main results. In Section 3, we review the specific construction of Wilton in [18] and bring in Jordan’s theorem. In Section 4, we introduce Markus-Epstein algorithm to provide a more concrete characterization of the structure of subgroups defined by labeled graphs. Finally, in Section 5, we employ Wilton’s construction and Markus-Epstein algorithm to prove Theorem 1.5.
2. Preliminaries
In this section, we review some notions and properties of LERFness and topology of graphs. Let $G=\langle X\mid R\rangle$ be a group and
[…]
2.1. Labeled graph and embedding
A graph is a 1-dimensional CW-complex consisting of two sets $\mathrm{V}$ and $\mathrm{E}$ representing the 0-cells and 1-cells respectively and two functions $\mathrm{E}\to\mathrm{E}$ and $\mathrm{E}\to\mathrm{V}$: for each $e\in\mathrm{E}$, there is an element $\bar{e}\in\mathrm{E}$ and an element $\iota(e)\in\mathrm{V}$. We call edge for $e\in\mathrm{E}$ and vertex for $v\in\mathrm{V}$, the requirements of $\bar{e}$ are $\bar{\bar{e}}=e$ and $\bar{e}\neq e$. A map of graph consists of a pair of functions preserving the structure, which maps edges to edges, vertices to vertices.
A graph $\Gamma$ is called a labeled graph (with the group $G=\langle X\mid R\rangle$) if it is equipped with a labeling function
$\mathrm{Lab}:\mathrm{E}(\Gamma)\to X^{\ast}$
satisfying $\mathrm{Lab}(\bar{e})=\mathrm{Lab}(e)^{-1}$ and the immersion condition: $e_{1}=e_{2}$ provided $\mathrm{Lab}(e_{1})=\mathrm{Lab}(e_{2})$ and $\iota(e_{1})=\iota(e_{2})$. The label of a path $p=e_{1}e_{2}\cdots e_{n}\subset\Gamma$ is defined by
$\mathrm{Lab}(p)\equiv\mathrm{Lab}(e_{1})\cdot\mathrm{Lab}(e_{2})\cdots\mathrm{Lab}(e_{n})\in G.$
If in addition, any path $p$ with $\mathrm{Lab}(p)=1$ is closed, then $\Gamma$ is called $G$-based. For each vertex $v$ and each word $g\equiv x_{1}x_{2}\cdots x_{n}$ , since $\Gamma$ is a labeled graph, there is a unique path $p$ in $\Gamma$ originating from $v$ with $\mathrm{Lab}(p)\equiv x_{1}x_{2}\cdots x_{n}$. The subgroup
$\mathrm{Lab}(\Gamma,v):=\{\mathrm{Lab}(p)\mid p~\textrm{is a loop in}~\Gamma\ \textrm{with base point}~v\}\leq G$
is called the subgroup of $G$ determined by the graph $\Gamma$.
Let $x\in X^{\ast}$. A vertex $v\in\mathrm{V}(\Gamma)$ is called $x$-saturated if there exists $e\in\mathrm{E}(\Gamma)$ with $\iota(e)=v$ and $\mathrm{Lab}(e)=x$. If $\Gamma$ is $x$-saturated for any $x\in X^{\ast}$ at any $v\in\mathrm{V}(\Gamma)$, then $\Gamma$ is called $X^{\ast}$-saturated.
[…]
Remark 2.1.
A labeled graph is also called a well-labeled graph in some literature, see [9] [10]. Any labeled graph of a free group $F_{r}$ is $F_{r}$-based.
$2.$ Relative Cayley graph
For a group $G=\langle X\mid R\rangle$, we denote the Cayley graph of $G$ with respect to the given generating set $X$ by $\mathrm{Cay}(G)$. Recall that $\mathrm{Cay}(G)$ is an oriented graph with the vertex set $\mathrm{V}(\mathrm{Cay}(G))=G$ and the edge set $\mathrm{\mathrm{E}}(\mathrm{Cay}(G))=G\times X^{\ast}$. An edge $(g,x)\in\mathrm{E}(\mathrm{Cay}(G))$ originates at $g$ and terminates at gxgx.
For any subgroup $H\leq G$, we denote by $\mathrm{Cay}(G,H)$ the Cayley graph of $G$ relative to the subgroup $H$ of $G$, which is an oriented graph with vertex set consisting of all the right cosets of $H$:
$\mathrm{V}(\mathrm{Cay}(G,H)):=\{Hg\mid g\in G\},$
and the edge set $\mathrm{E}(\mathrm{Cay}(G,H))=\mathrm{V}(\mathrm{Cay}(G,H))\times X^{\ast}$. We use $\mathrm{Cay}(G,H,H\cdot 1)$ when the base point $H\cdot 1$ is emphasized. An edge $(Hg,x)$ originates from the vertex HgHg and ends at HgxHgx. Moreover, a path $p$ originating from HgHg must end at $Hg\mathrm{Lab}(p)$. Therefore, we have the following.
Lemma $2.$
Let $H\leq G$ be a subgroup of $G$. A path $p$ originating from the base point $H\cdot 1$ in $\mathrm{Cay}(G,H,H\cdot 1)$ is a loop if and only if $\mathrm{Lab}(p)\in H$. Hence,
$\mathrm{Lab}(\mathrm{Cay}(G,H),H\cdot 1)=H.$
The following lemma shows that $G$-based graphs are subgraphs of the relative Cayley graphs.
Lemma 2.3.
[5, Lemma 1.5] Let $G=\langle X\mid R\rangle$ be a group and let $(\Gamma,v_{0})$ be a graph labeled with $X$. Denote $\mathrm{Lab}(\Gamma,v_{0})=H$. Then
  1. (1)
    $\Gamma$ is $G$-based if and only if it can be embedded in $\mathrm{Cay}(G,H,H\cdot 1)$;
  2. (2)
    $\Gamma$ is $G$-based and $X^{\ast}$-saturated if and only if it is isomorphic to $\mathrm{Cay}(G,H,H\cdot 1)$.
Note that $\mathrm{Cay}(G)$ and $\mathrm{Cay}(G,H)$ are naturally labeled with the generating set $X$ of $G$, and there are natural action of $G$ on $\mathrm{V}(\mathrm{Cay}(G))$ and $\mathrm{V}(\mathrm{Cay}(G,H))$ given by left translation. The relative Cayley graph $\mathrm{Cay}(G,H)$ can be defined as the quotient of $\mathrm{Cay}(G)$ by the action of $H$. Moreover, $\mathrm{Cay}(G,H)$ can be viewed as the 1-skeleton of a covering space of the representation complex of $G$, so we follow the terminology of Gitik [5] and refer to the "cover" as the relative Cayley graph.
Scott [13] described the residually finite and LERF properties by the embeddability from subgraphs to covers. Gitik [5] rephrased Scott’s result as follows.
Theorem 2.4.
[5, Theorem 1.1] Let $G=\langle X\mid R\rangle$ be a group.
  1. (1)
    $G$ is residually finite if and only if for any finite tree $\Gamma$, which is a subgraph of the Cayley graph of $G$, there exists a finite-index subgroup $H_{0}$ of G such that $\Gamma$ can be embedded in $\mathrm{Cay}(G,H_{0})$.
  2. (2)
    $G$ is LERF if and only if for any finitely generated subgroup $H$ of $G$ and for any finite connected subgraph $(\Gamma,H\cdot 1)$ of $\mathrm{Cay}(G,H,H\cdot 1)$, there exists a finite-index subgroup $H_{0}$ of $G$ such that $(\Gamma,H\cdot 1)$ can be embedded in $\mathrm{Cay}(G,H_{0},H_{0}\cdot 1)$.
2.3. Precover
The notion of precover was defined by Gitik in [5] and then actively employed by Markus-Epstein to study algorithmic problems in amalgams of finite groups in [9, 10].
Let $G_{1}$ and $G_{2}$ be two groups with generating sets $X$ and $Y$ respectively such that $X^{\ast}\cap Y^{\ast}=\emptyset$, and let $A=G_{1}\mathop{\ast}G_{2}$ be the free product of $G_{1}$ and $G_{2}$.
Definition 2.5.
Let $\Gamma$ be a graph labeled with $X\cup Y$.
  1. (1)
    A subgraph of $\Gamma$ is called monochromatic if it is labeled only with $X$ or only with $Y$. An $X$-component (resp. $Y$-component) of $\Gamma$ is a maximal connected subgraph of $\Gamma$ labeled with $X$ (resp. $Y$).
  2. (2)
    A vertex $v$ in $\mathrm{V}(\Gamma)$ is called $X$-monochromatic (resp. $Y$-monochromatic), if all edges connected to $v$ are labeled with $X$ (resp. $Y$). Otherwise, $v$ is called bichromatic.
  3. (3)
    $\Gamma$ is called a precover of AA, if it is AA-based, and each $X$-component of $\Gamma$ is a cover of $G_{1}$ and each $Y$-component of $\Gamma$ is a cover of $G_{2}$.
There is an equivalent description of the AA-based graph as follows.
Lemma 2.6 ([5]).
If $A=G_{1}\ast G_{2}$ is a free product, then any graph labeled with $X\cup Y$ is AA-based if and only if each $G_{i}$-component is $G_{i}$-based for $i=1,2$. In particular, it is AA-based if each monochromatic component is a cover of the corresponding free factor.
Lemma 2.7.
Let $\Gamma$ be a finite, connected, AA-based graph labeled with $X\cup Y$ and $\Gamma_{1}$ be an $X$-component. Let $P$ be a loop in $\Gamma$ starting from $v_{1}\in\Gamma_{1}$. If $1\neq\mathrm{Lab}(P)\in G_{1}$, then there is a loop $P^{\prime}$ in $\Gamma_{1}$ starting from $v_{1}$ such that $\mathrm{Lab}(P^{\prime})=\mathrm{Lab}(P)$.
Proof.
We denote $P=P_{1}P_{2}\cdots P_{n}$, each $P_{j}(1\leq j\leq n)$ is a maximal connected subpath labeled only with $X$ or $Y$. Then $\mathrm{Lab}(P_{j})\in G_{i}$ for some $i\in{1,2}$. Moreover, since
$\mathrm{Lab}(P)=\mathrm{Lab}(P_{1})\mathrm{Lab}(P_{2})\cdots\mathrm{Lab}(P_{n})\in G_{1}$
and AA is a free product of $G_{1}$ and $G_{2}$, we have $\mathrm{Lab}(P_{i})=1$ for each $P_{i}$ labeled with $Y$. Then $P_{i}$ must be closed because $\Gamma$ is AA-based. Moreover, as $P_{i}$ is maximal, both $P_{i-1}$ and $P_{i+1}$ are contained in $X$-components with
$\tau(P_{i-1})=\iota(P_{i})=\tau(P_{i})=\iota(P_{i+1}).$
It implies $P_{i-1}$ and $P_{i+1}$ are contained in the same $X$-component. So we can remove $P_{i}$ and combine $P_{i-1}$ with $P_{i+1}$ as a new monochromatic component. After removing all such $P_{i}$, we obtain a loop $P^{\prime}$ that starts from $v_{1}$ and is contained in the $X$-component $\Gamma_{1}$, such that $\mathrm{Lab}(P^{\prime})=\mathrm{Lab}(P)$. ∎
The precover of free products has a key property.
Lemma 2.8.
[5, Lemma 2.4] For any groups $G_{1}=\langle X\mid R\rangle$ and $G_{2}=\langle Y\mid T\rangle$, each precover of $G_{1}\ast G_{2}$ can be embedded in a cover of $G_{1}\ast G_{2}$ with the same set of vertices.
2.4. Pushout and folding
Pushout is an effective way to construct a new graph from old graphs considered by Stallings [14]. For example, in the category of labeled graphs, if $\alpha_{1}:\Gamma\to\Delta_{1}$ and $\alpha_{2}:\Gamma\to\Delta_{2}$ are injective maps, then the amalgam of $\Delta_{1}$ and $\Delta_{2}$ denoted by $\Delta_{1}\mathop{\ast}\limits_{\Gamma}\Delta_{2}$ is given by the pushout of $\alpha_{1}$ and $\alpha_{2}$:
$\Gamma$$\Delta_{1}$$\Delta_{2}$$\Delta_{1}\mathop{\ast}\limits_{\Gamma}\Delta_{2}$$\scriptstyle{\alpha_{1}}$$\scriptstyle{\alpha_{2}}$
In general, none of the graphs need be connected and none of the maps need be injective, moreover, pushouts do not always exist. The pushout of $\alpha_{1}:\Gamma\to\Delta_{1}$ and $\alpha_{2}:\Gamma\to\Delta_{2}$ exists if and only if there are orientations of $\Gamma$, $\Delta_{1}$ and $\Delta_{2}$ which are preserved by $\alpha_{1}$ and $\alpha_{2}$. We only consider the case that both $\alpha_{1}$ and $\alpha_{2}$ are injective maps, the amalgam can be constructed by performing the foldings.
Let $f:\mathrm{E}(\Gamma)\to X^{\ast}$ be a function on the graph $\Gamma$ such that $f(\bar{e})=(f(e))^{-1}$. For any two distinct edges $e_{1},e_{2}\in\mathrm{E}(\Gamma)$ with $\iota(e_{1})=\iota(e_{2})$ and $e_{1}\neq\bar{e}_{2}$, if $f(e_{1})=f(e_{2})$, then the folding $(e_{1},e_{2})$ is the projection of $\Gamma$ onto the quotient graph $\Gamma_{1}$ which is constructed by identifying $e_{1}$ with $e_{2}$ and $\tau(e_{1})$ with $\tau(e_{2})$. After performing a finite sequence of foldings as much as possible, the corresponding graph of $\Gamma$ is indeed a labeled graph. It can be easily seen that the above amalgam of $\alpha_{1}$ and $\alpha_{2}$ can be constructed by taking the disjoint union of graphs, identifying $\alpha_{1}(\Gamma)$ with $\alpha_{2}(\Gamma)$ and then performing finite foldings until a labeled graph is obtained.
3. Wilton’s proof of free groups
In this section, we recall the specific structure of Wilton’s proof of that any finitely generated infinite-index subgroup of a free group with rank greater than one is $\mathcal{A}$-separable.
3.1. Jordan’s theorem
An action of a group $H$ on $\{1,2,\ldots,n\}$ is called symmetric (resp. alternating) if the image of $H\to S_{n}$ is $S_{n}$ (resp. $A_{n}$). A subgroup $G\leq S_{n}$ is called transitive if it acts transitively on $\{1,2,\ldots,n\}$, and called primitive if in addition, it does not preserve any nontrivial partition.
The main tool for proving Lemma 3.3 is the following Jordan’s theorem, which gives a criterion for an action to be symmetric or alternating.
Theorem 3.1 (Jordan’s theorem, [4]).
For each $k>2$ there exists NN such that if $n>N$, $G\leq S_{n}$ is a primitive subgroup and there exists $\gamma\in G\setminus\{e\}$, which moves less than $k$ elements, then $G=S_{n}$ or $A_{n}$.
Note that if $G\leq S_{n}$ is transitive, then the cardinality of any partition preserved by $G$ divides $n$. Therefore, if $p$ is prime, then $G\leq S_{p}$ is primitive if and only if $G$ is transitive. We have an immediate corollary of Theorem 3.1 as follows.
Lemma 3.2.
For each $k>2$, there exists NN such that for any prime $p>N$ and any transitive subgroup $G\leq S_{p}$ if there exists $\gamma\in G\setminus\{e\}$, which moves less than $k$ elements, then $G=S_{p}$ or $A_{p}$.
3.2. Finite cover and Wilton’s construction
Let $F_{r}$ be a free group of rank $r>1$ with the generating set $X=\{x_{1},\cdots,x_{r}\}$. Stallings [14] showed: any finite graph labeled with $X$ can be embedded into a cover without increasing the number of vertices. Wilton made a modification in the construction to force the action of $F_{r}$ on the vertex set of the covering to be alternating or symmetric, see [18, Lemma 18 and Lemma 19]. We summarize them to the following lemma and describe Wilton’s specific construction in detail as a proof, which shall be used in our proof of Theorem 1.5.
Lemma $3.$
Let $Z$ be a finite, connected graph labeled with $X$. If $Z$ is not $X^{\ast}$-saturated, then there is a finite cover $Y$ of $F_{r}$ such that $Z$ embeds in $Y$ and the action of $F_{r}$ on $\mathrm{V}(Y)$ is alternating or symmetric.
Proof.
Suppose $Z$ has $k$ vertices. Then by Lemma 3.2, there exists a sufficiently large prime $p$ for $k+5$ such that for any transitive subgroup $G\leq S_{p}$ if there exists $\gamma\in G\setminus\{e\}$, which moves less than $k+5$ elements, then $G=S_{p}$ or $A_{p}$. We shall construct a finite cover $Y$ with $p$ vertices, such that $F_{r}$ acts on $\mathrm{V}(Y)$ transitively and there is an element of $F_{r}$ moving less than $k+5$ vertices.
Since $Z$ is not an $X^{\ast}$-saturated graph, there are (not necessarily distinct) vertices $a,b$ and some $x\in X$ such that $Z$ is not $x$-saturated at $a$ and not $x^{-1}$-saturated at $b$. Without loss of generality, we take $x=x_{1}$.
We begin by constructing two particular labeled graphs $W_{n}$ and $V_{s}$ with n=pΓkΓ4n=p-k-4 and 44 vertices respectively, such that $x_{i}\in X$ acts trivially on $\mathrm{V}(W_{n})$ for each $i\geq 2$ and $x_{2}$ acts nontrivially on $\mathrm{V}(V_{s})$.
$x_{1}$$x_{1}$$x_{1}$$x_{2}$$x_{2}$$x_{2}$$x_{2}$$w_{1}$$w_{2}$$w_{3}$$w_{4}$$v_{3}$$v_{4}$$v_{1}$$v_{2}$$x_{2}$$x_{2}$$x_{2}$$x_{1}$$x_{2}$$x_{1}$$x_{1}$
Figure 1. The graph $W_{4}$ on the left and $V_{+1,-1}$ on the right.
The construction of $W_{n}$ is as follows. We first take an interval with $n$ vertices, denoted by $w_{1},\ldots,w_{n}$ and denote the nΓ1n-1 edges by $e_{1},\ldots,e_{n-1}$. We label all edges with $x_{1}$ and orient these edges by
$\iota(e_{1})=w_{1},\quad\tau(e_{j})=\iota(e_{j+1})=w_{j+1},\quad\tau(e_{n-1})=w_{n}.$
Then for each generator $x_{i}\in X(i\geq 2)$ and each vertex $w_{j}(1\leq j\leq n)$, we attach an oriented edge labeled with $x_{i}$ from $w_{j}$ to itself. The resulting labeled graph is $W_{n}$. (See the left side of Figure 1.) For each $i\geq 2$, $x_{i}$ fixes all vertices of $W_{n}$, and hence acts trivially on $\mathrm{V}(W_{n})$. Note that the unsaturated vertices in $W_{n}$ are $w_{1}$ and $w_{n}$: $w_{1}$ is not $x_{1}^{-1}$-saturated, and $w_{n}$ is not $x_{1}$-saturated.
[…]
We now construct the desired $Y$ by attaching some edges to $Z\cup W_{n}\cup V_{s}$ as follows. First, for the six unsaturated vertices $a$, $b$ in $Z$, $w_{1}$, $w_{n}$ in $W_{n}$, and $v_{1}$, $v_{2}$ in $V_{s}$, we attach three edges labeled with $x_{1}$: one from $a$ to $w_{1}$, one from $w_{n}$ to $v_{1}$ and one from $v_{2}$ to $b$. The resulting graph $Y^{\prime}$ has $k+n+4=p$ vertices and it is a labeled graph. Finally, by adding some edges to $Y^{\prime}$, we obtain the finite cover $Y$ without adding new vertices:
[…]
Note that $Y$ is connected and the vertices $w_{1},\ldots,w_{n}$ in $Y$ are fixed by the action of $x_{2}$. It follows that $F_{r}$ acts on $\mathrm{V}(Y)$ transitively and $x_{2}$ moves less than pΓn+1=k+5p-n+1=k+5 vertices. Therefore, $Y$ meets the requirements specified in the first paragraph and the proof is complete. ∎
Remark 3.4.
Note that to enable connecting $W_{n}$ to $Z$, we require the existence of an unsaturated vertex on $Z$, meaning that $Z$ is not a cover of $F_{r}$. If there are multiple monochromatic components, we correspondingly require that a certain $F_{r}$-component is not a cover.
4. Kurosh decomposition structure derived from the precover
In this section, we introduce the Markus-Epstein algorithm (see [10]), which provides a method for constructing subgroup’s Kurosh decompositions via precovers. We employ this approach to characterize the structure of subgroups defined by labeled graphs.
Theorem 4.1 (Kurosh Subgroup Theorem, [7]).
Suppose $G=\ast_{i\in I}G_{i}$ is a free product and $H$ is a subgroup of $G$. Then $H\cong(\ast H_{j})\ast F$, where each $H_{j}$ is isomorphic to an intersection $H\cap G^{k_{i}}_{i}$ of $H$ with a conjugate of some $G_{i}$. Further, the set $\{H_{j}\}$ is unique up to conjugation and re-indexing, and the rank of the free group $F$ is uniquely determined.
Let $G=G_{1}\ast G_{2}$ be the free product of $G_{1}$ and $G_{2}$ with generating sets $X$ and $Y$ respectively. Let $\Gamma$ be a finite graph labeled with $X\cup Y$ and $v_{0}$ be the base point of $\Gamma$. Furthermore, we assume that $\Gamma$ is connected, $G$-based and denote
$\mathrm{Lab}(\Gamma,v_{0})=H\leq G.$
For a monochromatic component $C$ of $\Gamma$ with base point $v\in\mathrm{V}(C)$, let $T(C)$ be a spanning tree of $C$ with root vertex $v$. Since $\Gamma$ is connected, there is an approach path $P_{v}$ from $v_{0}$ to $v$, we denote
$\mathrm{Lab}(P_{v})\equiv g_{v}.$
Without loss of generality, we can assume that $P_{v}\cap C=\{v\}$. (Otherwise, since $C$ is connected, we can assume the subpath $P_{c}=P_{v}\cap C$ is connected, and then change $v$ to be the initial point of $P_{c}$ and remove $P_{c}$ from $P_{v}$.) If $v_{0}\in\mathrm{V}(C)$, we choose $v=v_{0}$. Let $\Gamma^{\prime}$ be a connected subgraph of $\Gamma$ obtained by removing edges in $\mathrm{E}(C)\setminus\mathrm{E}(T(C))$. More precisely, $\mathrm{V}(\Gamma^{\prime})=\mathrm{V}(\Gamma)$ and
$\mathrm{E}(\Gamma^{\prime})=(\mathrm{E}(\Gamma)\setminus\mathrm{E}(C))\cup\mathrm{E}(T(C)).$
Following the above assumption, Markus-Epstein’s algorithm is based on the following [10, Lemma 4.2].
Lemma 4.2.
$H=\langle g_{v}\mathrm{Lab}(C,v)g_{v}^{-1},\mathrm{Lab}(\Gamma^{\prime},v_{0})\rangle=g_{v}\mathrm{Lab}(C,v)g_{v}^{-1}\ast\mathrm{Lab}(\Gamma^{\prime},v_{0})$.
Lemma 4.2 shows that every monochromatic component $C$ with nontrivial $\mathrm{Lab}(C,v)$ corresponds to a free factor $g_{v}\mathrm{Lab}(C,v)g_{v}^{-1}$ of $\mathrm{Lab}(\Gamma,v_{0})$. Then we obtain the following free decomposition of $\mathrm{Lab}(\Gamma,v_{0})$.
Lemma 4.3.
Let $\Gamma$ be a finite, connected, $G$-based graph labeled with $X\cup Y$. If there are finitely many monochromatic components $C_{i}(1\leq i\leq k)$ which are not trees, then
$\mathrm{Lab}(\Gamma,v_{0})=\big(\ast_{1\leq i\leq k}\mathrm{Lab}(P_{i})\mathrm{Lab}(C_{i},v_{i})\mathrm{Lab}(P_{i})^{-1}\big)\ast\mathrm{Lab}(\Delta,v_{0}),$
where $P_{i}$ is an approach path from the base point $v_{0}$ of $\Gamma$ to the base point $v_{i}$ of $C_{i}$, the subgraph $\Delta$ is obtained from $\Gamma$ by removing all edges in each component $C_{i}$ that do not belong to its spanning tree $T(C_{i})$, and $\mathrm{Lab}(\Delta,v_{0})$ is a free group that does not conjugate into $G_{1}$ or $G_{2}$.
Proof.
Let $\mathrm{Lab}(P_{i})=g_{i}$. We proceed by induction on the number $k$ of monochromatic components which are not trees. Since there are finitely many such monochromatic components, the process is finite. When $k=1$, by Lemma 4.2,
$\mathrm{Lab}(\Gamma,v_{0})=g_{1}\mathrm{Lab}(C_{1},v_{1})g_{1}^{-1}\ast\mathrm{Lab}(\Gamma^{\prime},v_{0}),$
$\Gamma^{\prime}$ is the required $\Delta$.
When $k>1$, we start from the monochromatic component $C_{1}$ containing the base point $v_{0}$ of $\Gamma$ such that $v_{1}=v_{0}\in\mathrm{V}(C_{1})$, then by Lemma 4.2,
$\mathrm{Lab}(\Gamma,v_{0})=g_{1}\mathrm{Lab}(C_{1},v_{1})g_{1}^{-1}\ast\mathrm{Lab}(\Gamma^{\prime\prime},v_{0}).$
Since $\Gamma^{\prime\prime}$ embeds in $\Gamma$ and has fewer monochromatic components which are not trees than $\Gamma$, by the induction hypothesis, we have
$\mathrm{Lab}(\Gamma^{\prime\prime},v_{0})=(\ast_{2\leq i\leq k}~g_{i}\mathrm{Lab}(C_{i},v_{i})g_{i}^{-1})\ast\mathrm{Lab}(\Delta,v_{0}).$
By [10, Lemma 5.2], $\mathrm{Lab}(\Delta,v_{0})$ is a free group that does not conjugate into $G_{1}$ or $G_{2}$. Then combining the above two equation, the proof is complete. ∎
Moreover, the free decomposition in Lemma 4.3 coincides with the standard Kurosh decomposition.
Lemma $4.$
In Lemma 4.3, each free factor
$L_{i}:=\mathrm{Lab}(P_{i})\mathrm{Lab}(C_{i},v_{i})\mathrm{Lab}(P_{i})^{-1}=\mathrm{Lab}(\Gamma,v_{0})\cap G_{k}^{\mathrm{Lab}(P_{i})}$
for some $k\in\{1,2\}$ and $\mathrm{Lab}(C_{i},v_{i})\leq G_{k}$.
Proof.
Since $C_{i}$ is a monochromatic component, $\mathrm{Lab}(C_{i},v_{i})\leq G_{k}$ for some $k\in\{1,2\}$ and
$L_{i}\leq\mathrm{Lab}(\Gamma,v_{0})\cap G_{k}^{\mathrm{Lab}(P_{i})}$
clearly. Below we show $\mathrm{Lab}(\Gamma,v_{0})\cap G_{k}^{\mathrm{Lab}(P_{i})}\leq L_{i}.$
For any $\gamma\in\mathrm{Lab}(\Gamma,v_{0})\cap G_{k}^{\mathrm{Lab}(P_{i})}$, we can denote $\gamma=\mathrm{Lab}(P_{i})g\mathrm{Lab}(P_{i})^{-1}$ for some $g\in G_{k}$. Since $\gamma\in\mathrm{Lab}(\Gamma,v_{0})$, there is a loop $P_{\gamma}$ starting from $v_{0}$ with $\gamma=\mathrm{Lab}(P_{\gamma})$. It implies that
$g=\mathrm{Lab}(P_{i})^{-1}\mathrm{Lab}(P_{\gamma})\mathrm{Lab}(P_{i})=\mathrm{Lab}(\bar{P}_{i}P_{\gamma}P_{i})\in G_{k},$
where $\bar{P}_{i}P_{\gamma}P_{i}$ is a loop starting from $v_{i}$. By Lemma 2.7, there is a loop $P_{g}\subset C_{i}$ starting from $v_{i}$ such that
$g=\mathrm{Lab}(\bar{P}_{i}P_{\gamma}P_{i})=\mathrm{Lab}(P_{g}),$
which implies that $g\in\mathrm{Lab}(C_{i},v_{i})$ and then $\gamma\in L_{i}$. ∎
5. Proof of the main theorem
In this section, we employ Wilton’s construction and Markus-Epstein’s algorithm to establish Theorem 1.5.
[…]
$H=\langle h_{1},h_{2},\cdots,h_{n}\rangle\leq F_{r}\ast G.$
For any finite set $\{\gamma_{1},\cdots,\gamma_{m}\}\subset(F_{r}\ast G)\setminus H$, we pick a finite connected subgraph $\Gamma$ of $\mathrm{Cay}(F_{r}\ast G,H)$ as follows:
($\ast$) $\Gamma=\bigcup_{i=1}^{n}p_{i}\bigcup_{j=1}^{m}p^{\prime}_{j}\subset\mathrm{Cay}(F_{r}\ast G,H),$
where all the loops $p_{i}$ and paths $p^{\prime}_{j}$ have the same initial point $H\cdot 1$. and the labels of $p_{i}$ and $p^{\prime}_{j}$ are given by $h_{i}$ and $\gamma_{j}$ respectively. Note that $\Gamma$ is an $F_{r}\ast G$-based labeled graph because it embeds in the $F_{r}\ast G$-based graph $\mathrm{Cay}(F_{r}\ast G,H)$. Moreover, any $F_{r}$-component of $\Gamma$ is $F_{r}$-based and any $G$-component of $\Gamma$ is $G$-based by Lemma 2.6.
Lemma 5.1.
$\mathrm{Lab}(\Gamma,H\cdot 1)=\mathrm{Lab}(\mathrm{Cay}(F_{r}\ast G,H),H\cdot 1)=H$.
Proof.
The last “=" follows from Lemma $2.$ Since $\Gamma$ is the subgraph of $\mathrm{Cay}(F_{r}\ast G,H)$, each loop in $\mathrm{Lab}(\Gamma,H\cdot 1)$ is again a loop in $\mathrm{Lab}(\mathrm{Cay}(F_{r}\ast G,H),H\cdot 1)$, that is,
$\mathrm{Lab}(\Gamma,H\cdot 1)\leq\mathrm{Lab}(\mathrm{Cay}(F_{r}\ast G,H),H\cdot 1).$
Conversely, since the labels of the loops $p_{i}(1\leq i\leq n)$ are given by the generators $h_{i}$ of $H$, we have $H\leq\mathrm{Lab}(\Gamma,H\cdot 1)$ clearly. ∎
We can describe the $\mathcal{A}\cup\mathcal{S}$-separability of $H$ via the embedding property of $\Gamma$.
Proposition 5.2.
[…]
Proof.
Following the notation above, we will show that there exists a surjection $f$ from $F_{r}\ast G$ to an alternating or symmetric group such that $f(\gamma_{j})\notin f(H)$ for all $1\leq j\leq m$.
By the hypothesis, there exists a subgroup $H_{0}\leq F_{r}\ast G$ with prime index $p$ as in Lemma 3.2 such that $(\Gamma,H\cdot 1)$ embeds in $\mathrm{Cay}(F_{r}\ast G,H_{0},H_{0}\cdot 1)$. Note that
[…]
and the action of $F_{r}\ast G$ on $\mathrm{V}(\mathrm{Cay}(F_{r}\ast G,H_{0}))$ is transitive and there exists $\gamma\in F_{r}\ast G$ moves less than $k$ vertices, then we have a homomorphism
$f:F_{r}\ast G\to S_{p}$
with $f(F_{r}\ast G)=S_{p}$ or $A_{p}$ by Lemma 3.2. Moreover, since for any $j=1,\ldots,m$,
$\gamma_{j}\notin H=\mathrm{Lab}(\mathrm{Cay}(F_{r}\ast G,H),H\cdot 1),$
the path $p^{\prime}_{j}$ is not closed in $\Gamma$ and hence not closed again in $\mathrm{Cay}(F_{r}\ast G,H_{0})$. Therefore, the action of $\gamma_{j}$ on $\mathrm{V}(\mathrm{Cay}(F_{r}\ast G,H_{0},H_{0}\cdot 1))$ does not fix the base point $H_{0}\cdot 1$ while the action of $H$ does, it implies $f(\gamma_{j})\notin f(H)$ for any $1\leq j\leq m$. ∎
Lemma 5.3.
If $H\cap F_{r}^{\gamma}\neq 1$ has infinite index in $F_{r}^{\gamma}$ for some $\gamma\in F_{r}\ast G$, then there is an $F_{r}$-component in $\Gamma$ which is not a cover of $F_{r}$.
Proof.
By Lemma 5.1, we have $H=\mathrm{Lab}(\Gamma,v_{0})$ with base point $v_{0}:=H\cdot 1$. Then combining Kurosh subgroup theorem with Lemma 4.4, there is an $X$-component $C$ of $\Gamma$ such that
$\gamma\mathrm{Lab}(C,v)\gamma^{-1}=H\cap F_{r}^{\gamma},$
where $\gamma=\mathrm{Lab}(P)$ for $P$ an approach path from the base point $v_{0}$ to the base point $v$ of $C$. Hence, the index
$F_{r}:\mathrm{Lab}(C,v)]=[F_{r}^{\gamma}:\gamma\mathrm{Lab}(C,v)\gamma^{-1}]=[F_{r}^{\gamma}:H\cap F_{r}^{\gamma}]=\infty.$
It follows that the $X$-component $C$ is not a finite cover of $F_{r}$. In fact, $C$ is not a cover of $F_{r}$ because it is finite. ∎
We now proceed to prove Theorem 1.5.
Theorem 1.5.
Let $G$ be a LERF group and let $H\leq F_{r}\ast G$ be a finitely generated subgroup. Then $H$ is $\mathcal{A}\cup\mathcal{S}$-separable in $F_{r}\ast G$ if one of the following holds:
  1. (1)
    $H\cap F_{r}^{\gamma}=1$ for every $\gamma\in F_{r}\ast G$, or
  2. (2)
    $H\cap F_{r}^{\gamma}\neq 1$ has infinite index in $F_{r}^{\gamma}$ for some $\gamma\in F_{r}\ast G$.
Proof.
By Proposition 5.2, we need to find a subgroup $H_{0}\leq F_{r}\ast G$, or equivalently, a cover
$\widetilde{\Gamma}:=\mathrm{Cay}(F_{r}\ast G,H_{0},H_{0}\cdot 1),$
[…]
We construct $\widetilde{\Gamma}$ in the following four steps.
Step 1. Embedding of each monochromatic component into a finite cover.
Since $\Gamma$ is finite, there must be finitely many monochromatic components in $\Gamma$, we denote them by $\Gamma_{1},\ldots,\Gamma_{\ell}$, respectively.
If hypothesis (1) holds, either there is no $F_{r}$-component in $\Gamma$ (i.e., $\Gamma$ is $Y$-monochromatic), or every $F_{r}$-component in $\Gamma$ is a finite tree. Then for the former case, we denote the base point $H\cdot 1$ by $\Gamma_{1}$ and re-index the only $Y$-monochromatic component as $\Gamma_{2}$. For the latter case, we assume such an $F_{r}$-component is $\Gamma_{1}$ which is not a cover of $F_{r}$.
If hypothesis (2) holds, then by Lemma 5.3, there is an $F_{r}$-component in $\Gamma$ which is not a cover of $F_{r}$. Without loss of generality, we again assume such an $F_{r}$-component is $\Gamma_{1}$.
Recall that each finite connected labeled component $\Gamma_{i}$ is $F_{r}$- or $G$-based, then $\Gamma_{i}$ can be embedded in a cover of $F_{r}$ or $G$ by Lemma 2.3. Moreover, since both $F_{r}$ and $G$ are LERF, every $\Gamma_{i}$ can be embedded in a finite cover $\widetilde{\Gamma}_{i}$ of $F_{r}$ or $G$ respectively by Theorem 2.4.
Step 2. Construction of $\Gamma^{\ast}$.
We now consider the pushout of the embedding $\alpha_{2}:\bigcup_{i\geq 2}\Gamma_{i}\hookrightarrow\bigcup_{i\geq 2}\widetilde{\Gamma}_{i}$ as follows:
$\mathop{\bigcup}\limits_{i\geq 2}\Gamma_{i}$$\Gamma$$\mathop{\bigcup}\limits_{i\geq 2}\widetilde{\Gamma}_{i}$$\Gamma^{\ast}$$\scriptstyle{\alpha_{1}}$$\scriptstyle{\alpha_{2}}$
Note that every $\Gamma_{i}$ embeds in $\widetilde{\Gamma}_{i}$ and $\Gamma_{i}$ is a monochromatic component of $\Gamma$, so there are no foldings occur between distinct $\widetilde{\Gamma}_{i}$, it follows $\Gamma_{i}$ embeds in $\Gamma^{\ast}$. Moreover, every monochromatic component of $\Gamma^{\ast}$ is $F_{r}$- or $G$-based, hence $\Gamma^{\ast}$ is $F_{r}\ast G$-based by Lemma 2.6. (See Figure 2.)
$y_{1}$$y_{1}$$Hy_{1}^{-1}$$Hy_{1}$$x_{1}$$Hy_{2}$$H$$Hx_{1}$$x_{1}$$x_{2}$$y_{2}$$x_{1}$$y_{1}$$y_{1}$$y_{1}$$y_{2}$$y_{2}$$x_{1}$$x_{1}$$x_{2}$$y_{1}$$y_{2}$$y_{2}$$x_{2}$$y_{2}HH$$Hx_{1}$$Hy_{1}^{-1}$$Hy_{1}$
Figure 2. An example for $\Gamma$ and $\Gamma^{\ast}$ when $X=\{x_{1},x_{2}\}$, $Y=\{y_{1},y_{2}\}$. We take $H=\langle y_{1}x_{1}^{-1}y_{1},x_{1}x_{2}x_{1}^{-1}\rangle$ and $\gamma_{1}=y_{2}$.
Step 3. Modify $\Gamma_{1}$ to obtain a precover $\Gamma^{\prime}$ of $F_{r}\ast G$.
[…]
We now take a sufficiently large prime $p$ for $k+5$ as in Lemma 3.2, and let n=pΓkΓ4n=p-k-4. We construct the pushout of the embedding $\alpha_{2}^{\prime}:\Gamma_{1}\hookrightarrow\widetilde{\Gamma}_{1}$ as following:
$\Gamma_{1}$$\Gamma^{\ast}$$\widetilde{\Gamma}_{1}$$\Gamma^{\prime}$$\scriptstyle{\alpha^{\prime}_{1}}$$\scriptstyle{\alpha^{\prime}_{2}}$
Notice that $\Gamma_{1}$ embeds in $\widetilde{\Gamma}_{1}$ and $\widetilde{\Gamma}_{1}$ is monochromatic, there are no foldings occur between $\widetilde{\Gamma}_{1}$ and $\Gamma^{\ast}$, then both $\Gamma^{\ast}$ and $\widetilde{\Gamma}_{1}$ embed in $\Gamma^{\prime}$. Since $\Gamma^{\ast}$ is $F_{r}\ast G$-based and $\widetilde{\Gamma}_{1}$ is $F_{r}$-based, $\Gamma^{\prime}$ is $F_{r}\ast G$-based by Lemma 2.6 and every monochromatic component is a cover. So $\Gamma^{\prime}$ is a precover of $F_{r}\ast G$.
Step 4. Construction of the final cover $\widetilde{\Gamma}$.
By Lemma 2.8, the precover $\Gamma^{\prime}$ can embed in a cover $\widetilde{\Gamma}$ of $F_{r}\ast G$ without adding new vertices, so $\Gamma$ itself embeds in the cover $\widetilde{\Gamma}$ with
[…]
Since $x_{2}$ acts trivially on $\mathrm{V}(W_{n})$ and acts nontrivially on $\mathrm{V}(V_{s})$, there must be less than $k+5$ vertices (lying in $\Gamma^{\ast}\cup V_{s}$ ) in $\widetilde{\Gamma}$ moved by $x_{2}$. Therefore, the cover $\widetilde{\Gamma}$ is precisely what we need, and the proof is complete. ∎
Acknowledgements. The authors would like to thank Qian Chen very much for his comments. Thanks are also extended to the editor and the anonymous referee for their time.